2. Now, go back to the original supply and demand equations: Ed 60 2*w Es =-10+5w Redraw these on a new graph, as you did in the above question Suppose that a minimum wage of S12/hr is now imposed. a) What is the quantity demanded at this new wage? What is the quantity supplied? How many workers are hired? How many jobs are lost? b) Show all of the above on your graph. Once again, with the minimum wage in place, suppose that the demand again changes to: c) Repeat part a) d) Repeat part b).

Consider the same demand curve for labor as in the above questions: Ed 60 2*w Using the equilibrium wage and employment from the above question, calculate the 3. a) elasticity of demand when the $12/hr minimum wage is imposed. Suppose that the demand curve was now Ed 80-4*w b) Verify that the equilibrium wage and employment levels are the same. c) Calculate the elasticity of demand for this new demand curve when the minimum wage is imposed. d) I hich of these would you expect the employment effect of the minimum wage to be larger? Briefly explain.
Answer #1

2

a) Ed= 60-2*w

Es= -10+5*w

Minimum wage rate: $12/hr

Ed= 60-2*12

= 60-24

Ed= 36

Es= -10+5*w

= -10+5*12

= -10+60

Es= 50

Quantity of labor demanded= 36

Quantity of labor supplied= 50

36 workers are hired as per demand for the labor at wage rate $12/hr and 50 workers are supplied at wage rate $12/hr

b) Graph is as per below.

Labes

In the above graph it is shown that at wage rate $12/hr, Quantity of labor demanded is 36 and quantity of labor supplied is 50 which is depicted on demand and supply of labor curve and here DL is the Demand for Labor curve and SL curve is the supply of labor.

This shows supply of labor is in excess of demand of labor.

Now demand change to as per below

Ed' = 67-2*w

wage rate= $12/hr

Ed'= 67-2*12

Ed'= 67-24

Ed'= 43

c) Ed'= 43

Es= 50(It remains same as before)

Quantity of labor demanded= 43

Quantity of labor supplied= 50

d) As per above calculation new graph is as per below.

De mana and up 업 labeY LabesAs per above graph with revised demand, quantity of labor demanded is 43 and quantity of labor supplied is still 50

This shows supply of labor is in excess in comparison of demand for labor.